OpenStudy (anonymous):

given the table of M&M distributions below , P(not blue), the probability of selecting anything but blue M&M at random from a large bag RED 0.13 ORANGE 0.20 YELLOW ? GREEN 0.16 BLUE 0.24 BROWN 0.13

OpenStudy (campbell_st):

Well you can do this by using complementary events P(not blue) = 1 - P(blue)

OpenStudy (anonymous):

HOW @campbell_st

OpenStudy (campbell_st):

well you are given the probability of selecting a blue M & M P(blue) = 0.24 the maximum probability is always 1 so you have 2 options for finding the probability of not blue 1. add up the probabilities of all the other colours 2. use complementary events P(not blue) = 1 - P(blue) = 1 - 0.24 you'll get the same answer with either method..

OpenStudy (anonymous):

SOO I JUST DO THIS 1-0.24=0.76 @campbell_st

OpenStudy (campbell_st):

that's it when I do the not probability questions I always use P(not ?) = 1 - P(?) its just easier